You have graduated from college but unfortunately have $39,000 in outstanding loans. The loans require monthly payments of $268.33 (or $3,220 per year), which covers interest and principal repayment (that is, the loan has the same basic features as a mortgage). If the interest rate is 3 percent, how long will it take you to repay the debt? Use Appendix D to answer the question. Round your answer up to the next whole number.

If the powers that be raise the rate to 5 percent, how many additional years will be required to retire the loans? Use Appendix D to answer the question. Round your answer up to the next whole number.
Answer #1

Calculating Time Period Required to repay,

Using TVM Calculation,

N = [PV = 39,000, PMT = -268.33, FV = 0, I = 0.03/12]

N = 181 months

Interest Rate = 5%

Using TVM Calculation,

N = [PV = 39,000, PMT = -268.33, FV = 0, I = 0.05/12]

N = 224 months

Additional Time Period Required = (224 - 181)/12

Additional Time Period Required = 3.58 years